### v2

@@ 4 Simulations and Discussion
All code that was used to generate the figures in this section can be found at the GitHub repository here: https://github.com/kdayers/NoisyFixedPoints. Suppose $X$ is a random variable taking values in $[0,1]$. If $Pr\{X\leq x\}=Pr\{f(X)\leq x\}$ then we say that the distribution of $X$ is invariant under $f$. In this way, we can discuss invariant distributions for the deterministic logistic map, by looking at distribtuions invariant under $f_{\lambda}(x)=\lambda x(1-x)$. In [20], Ulam and von Neumann proved that the continuous invariant distribution for the deterministic logistic map with $\lambda=4$ is a beta distribution with $\alpha=\beta=0.5$. This function is symmetric on the interval (0,1), and goes to infinity at the endpoints of its support. Thus, the invariant distribution when $\lambda=4$ is given by the function
@@
Figure 1: Histogram of the beta distribution with $\alpha=\beta=0.5$. The solid line shows the theoretical distribution given in Equation (6); the bars are a histogram generated by sampling 100,000 initial values uniformly on $(0,1)$
However, invariant densities for other values of $\lambda$ are not, in general, symmetric. In [21], Hall and Wolff provided numerical estimates for the invariant distributions for the deterministic logistic map for various values of $\lambda$. They demonstrated that any continuous invariant density will have poles (singularities) at the boundary of its support, as well as additional poles at the interior of the support in the case $\lambda<4$. They also show that these invariant distributions were not symmetric. Figure 2 shows numerical approximations to theses deterministic invariant densities for $\lambda=3.87$ and $\lambda=3.95$ These numerical approximations are histograms of 100,000 sampled initial values iterated 2,000 times. They indicate the singularities in the theoretical densities described in Remark 5 of [21]. These singularities, or poles, correspond to the orbit of the critical point, $f_{\lambda}(1/2)$. In addition, Hall and Wolff observed that the invariant density will be supported on the interval $(\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda)$. This is due to the fact that any point in the interval $(\frac{1}{4}\lambda,1)$ will be mapped into the interval $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}))$, and any point in $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})$ will be mapped to a point in $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$. Therefore, after finitely many iterations, any point $x\in(0,1)$ must be mapped into $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$, and so the invariant density must be supported on this subinterval. Note that as $\lambda\rightarrow 4$, $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})\rightarrow 0$ and $\frac{1}{4}\lambda\rightarrow 1$, so the interval of support widens as $\lambda$ approaches 4.
@@
In each case, after 2,000 iterations, we obtained a histogram very similar to that presented in Figure 4: a bimodal, asymmetric distribution that appears to be supported throughout the unit interval, without singularities (poles) in the interior. This is noteworthy, particularly because we simulated using both continuous and discrete starting distributions. We examined distributions that were left-skewed, right-skewed, and symmetric. Each simulated initial distribution appears to converge to the same distribution after 2,000 iterations, aligning with the global asymptotic stability of the invariant distribution established theoretically earlier in this paper. We also note that, while the distribution demonstrates peaks toward the sides of the interval, it does not exhibit the poles at the boundary of the interval in the same manner as the invariant distributions for the deterministic maps. However, this seems reasonable, as the interval of support of the invariant distribution is dependent on $\lambda$, as discussed above, and relatively fewer values of $\lambda$ will map points into ranges closer to 0 and 1, so it makes sense that the distribution will decrease near the boundary of $[0,1]$. It is also of note that even as the width of the interval that the distribution of $\lambda$ is supported on narrows, the invariant distribution maintains this property of decreasing near the boundary of support. For instance, we ran simulations where $\lambda$ was sampled uniformly on the interval $(3.87,9)$ and again on the interval $(3.87,3.935).$ Both invariant distributions exhibited this decrease toward the edge of the boundary of support (which, as expected, is not the entirety of $(0,1)$. Please see Figure 5.
As for the speed of convergence, it appears that the convergence happens relatively quickly. Figure 6 shows the evolution of the simulated invariant density after various numbers of iterations, starting from a uniform distribution.
@@
Figure 4: Empirical distribution of $\{X_{n}\}$ after $n=2,000$ iterations, where 100,000 initial values were sampled from a truncated exponential distribution with parameter $r=1.25$ as in Figure 3. Similar results were obtained using other initial distributions (results not shown).
@@
Figure 6: The densities for a simulation of 100,000 values, initially sampled randomly from a uniform distribution on (0,1), after various numbers of iterations.
@@ 5 Conclusion
In conclusion, we have illustrated that the stochastic logistic map, characterized by parameter values uniformly distributed in the chaotic regime, exhibits Harris irreducibility. Consequently, the previously established invariant distribution stands as the unique invariant distribution. This result also implies that all initial distributions converge to the invariant distribution in the Cesàro sense. Moreover, we have demonstrated the strong aperiodicity of the stochastic logistic map, indicating that initial distributions converge to the invariant distribution with respect to the total variation distance—a more robust and rigorous result. We conducted simulations of the stochastic logistic map to visualize the invariant distribution. Remarkably, the empirically derived invariant distribution exhibits qualitative distinctions when compared to the invariant distributions of the deterministic logistic map for values of $\lambda$ in the chaotic regime. In particular, the simulated distribution does not appear to have a pole at $x=0$, which contrasts with the work in [21], which demonstrates that the deterministic invariant density will have a pole at the boundary of its support.

### v3

@@ 4 Simulations and Discussion
All code that was used to generate the figures in this section can be found at the GitHub repository here: https://github.com/kdayers/NoisyFixedPoints. Suppose $X$ is a random variable taking values in $[0,1]$. If, for all $x$, $Pr\{X\leq x\}=Pr\{f(X)\leq x\}$ then we say that the distribution of $X$ is invariant under $f$. In this way, we can discuss invariant distributions for the deterministic logistic map by looking at distributions invariant under $f_{\lambda}(x)=\lambda x(1-x)$. In [21], Ulam and von Neumann proved that the continuous invariant distribution for the deterministic logistic map with $\lambda=4$ is a beta distribution with $\alpha=\beta=0.5$. The density for this distribution is symmetric on the interval (0,1), and goes to infinity at the endpoints of its support. To be specific, the invariant distribution when $\lambda=4$ is given by the density function
@@
Figure 1: Histogram of the beta distribution with $\alpha=\beta=0.5$. The solid line shows the theoretical distribution given in Equation (6); the bars are a histogram generated by sampling 100,000 initial values uniformly on $(0,1)$ and then iterating 20 times.
However, invariant densities for other values of $\lambda$ are not, in general, symmetric. In [22], Hall and Wolff provided numerical estimates for the invariant distributions for the deterministic logistic map for various values of $\lambda$. They demonstrated that any continuous invariant density will have poles (singularities) at the boundary of its support, as well as additional poles at the interior of the support in the case $\lambda<4$. They also show that these invariant distributions were not symmetric. Figure 2 shows numerical approximations to theses deterministic invariant densities for $\lambda=3.87$ and $\lambda=3.95$ These numerical approximations are histograms of 100,000 sampled initial values iterated 20 times. They indicate the singularities in the theoretical densities described in Remark 5 of [22]. These singularities, or poles, correspond to the orbit of the critical point, $f_{\lambda}(1/2)$. In addition, Hall and Wolff observed that the invariant density will be supported on the interval $(\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda)$. This is due to the fact that any point in the interval $(\frac{1}{4}\lambda,1)$ will be mapped into the interval $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}))$, and any point in $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})$ will be mapped to a point in $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$. Therefore, after finitely many iterations, any point $x\in(0,1)$ must be mapped into $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$, and so the invariant density must be supported on this subinterval. Note that as $\lambda\rightarrow 4$, $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})\rightarrow 0$ and $\frac{1}{4}\lambda\rightarrow 1$, so the interval of support widens as $\lambda$ approaches 4.
@@
In each case, after several iterations, we obtained a histogram very similar to that presented in Figure 4: a bimodal, asymmetric distribution that appears to be supported throughout the unit interval, without singularities (poles) in the interior. This is noteworthy, particularly because we simulated using both continuous and discrete initial distributions. We examined distributions that were left-skewed, right-skewed, and symmetric. Each simulated initial distribution appears to converge to the same distribution, aligning with the global asymptotic stability of the invariant distribution established theoretically earlier in this paper. We also note that, while the distribution demonstrates peaks toward the sides of the interval, it does not exhibit the poles at the boundary of the interval in the same manner as the invariant distributions for the deterministic maps. However, this seems reasonable, as the interval of support of the invariant distribution is dependent on $\lambda$, as discussed above, and relatively fewer values of $\lambda$ will map points into ranges closer to 0 and 1, so it makes sense that the distribution will decrease near the boundary of $[0,1]$. It is also of note that even as the width of the interval on which $\lambda$ is supported narrows, the invariant distribution maintains this property of decreasing near the boundary of support. For instance, we ran simulations where $\lambda$ was sampled uniformly on the interval $(3.87,9)$ and again on the interval $(3.87,3.935).$ Both invariant distributions exhibited this decrease toward the edge of the boundary of support (which, as expected, is not the entirety of $(0,1)$. Please see Figure 5.
As for the speed of convergence, it appears that the convergence happens relatively quickly. Figure 6 shows the evolution of the simulated invariant density after various numbers of iterations, starting from a uniform distribution. In this set of simulations, the empirical distributions after 5, 20 and 200 iterations are quite close to each other, and the distribution after 20 iterations is almost indistinguishable from the distribution after 200 iterations. This justifies our use of only 20 iterates in the other simulations presented here.
@@
Figure 4: Empirical distribution of $\{X_{n}\}$ after $n=20$ iterations, where 100,000 initial values were sampled from a truncated exponential distribution with parameter $r=1.25$ as in Figure 3. Similar results were obtained using other initial distributions (results not shown).
@@
Figure 6: The densities for a simulation of 100,000 values, initially sampled randomly from a uniform distribution on (0,1), after various numbers of iterations where $\lambda$ is uniformly randomly sampled from $(3.87,4)$ each iteration.
@@ 5 Conclusion
In conclusion, we have illustrated that the stochastic logistic map, characterized by parameter values uniformly distributed in the chaotic regime, is irreducible. Consequently, the previously established invariant distribution stands as the unique invariant distribution. This result also implies that all initial distributions converge to the invariant distribution in the Cesàro sense. Moreover, we have demonstrated the strong aperiodicity of the stochastic logistic map, indicating that initial distributions converge to the invariant distribution with respect to the total variation distance—a more robust result. We conducted simulations of the stochastic logistic map to visualize the invariant distribution. Remarkably, the empirically derived invariant distribution exhibits qualitative distinctions when compared to the invariant distributions of the deterministic logistic map for values of $\lambda$ in the chaotic regime. In particular, the simulated distribution does not appear to have a pole at $x=0$. This contrasts with the work in [22], which demonstrates that the deterministic invariant density will have a pole at the boundary of its support.